OpenStudy (anonymous):

Can someone help me with this question? Find all values of a, for which the equation: square root of(a*x^2+ax+2)=ax+2 has a unique root.

OpenStudy (anonymous):

a*x^2+ax+2=ax+2 ax^2 = 0 x^2 = 0 x = 0 no matter what the value of a is, the equation always has a unique solution. Namely 0

OpenStudy (anonymous):

but there is a square root so I would get ax^2+ax+2=(ax+2)^2

OpenStudy (anonymous):

didn't see the square root -.- pft.. i'm off

OpenStudy (anonymous):

After some simplification, the first line below will transform to the second line \[ (a x^2+a x+2)=(a x+2)^2 \\ (a - a^2) x^2-3 a x -2 =0 \] One a=1 will make it a first degree equation with one root. a=0 does not. There is still another a. Can you find it?

OpenStudy (anonymous):

The discriminant of the quadratic equation above is \[ (-3 a)^2 -4(a-a^2)(-2)=8a +a^2= a(8+a)=0 a=0\\ a=-8 \] a=0 is not good. a =-8 will give you double root.

OpenStudy (anonymous):

for a=-8 the quadratic becomes \[ -2 + 24 x - 72 x^2 = 0\\ 36 x^2 - 12 x +1=0\\ (6x-1)^2=0\\ x= \frac 1 6 \]